# Challenges and recommendations for Electronic Health Records data extraction and preparation for dynamic prediction modelling in hospitalized patients - a practical guide

### v1

@@ 4.4 Data cleaning
Outcomes documented in the system are subject to mislabelling and delay in registration Assessments done during the patient’s stay, e.g. delirium assessment by a neurologist using Confusion Assessment Method (CAM) or Delirium Rating Scale (DRS), can be used to define the outcome. These can though be subject of different documentation practices in different wards. For example delirium might be consistently evaluated on the neurology ward and after surgery, but inconsistently in other wards or in patients that did not undergo surgery. Unless meticulously documented, clinical assessments are subject to inconsistencies in different patient groups or across medical wards. There may also be delays in documentation compared to the clinical assessment time (e.g.: documented at the end of the day), which can impact time-sensitive outcomes. Understanding the assessment practices of the event(s) identifying the outcome is essential in defining its reliability. Understanding the documentation practices (and the lags compared to the actual event time) can also be important, depending on how time-sensitive the prediction are, how often the predictions are renewed and the prediction time horizon. This information is typically retrieved from clinical experts, as documenting every single clinical assessment practice during data extraction would be overwhelming. Manual relabelling by clinical experts post extraction, based on patient assessments, possibly combined with other data such as clinical notes, can prove as a viable solution for some outcomes, although it remains a labour intensive task. Informativeness - understandability; Measurement process - human-introduced error - noisy labels; Measurement process - completeness; Timeliness - currency Correctness; Currency
Outcomes derived using clinical or surveillance definitions often require extraction of multiple data sources Clinical or surveillance definitions provide a more robust alternative compared to ICD codes or documentation of clinical assessments for certain acute events (e.g.: AKI, CLABSI, VAE). It comes with advantages: the event occurrence/onset timestamp can be established and the same definition is used over the entire training and test data, but its consistency might still depend on the recording practices of the items used in the definition. Furthermore, these items span across multiple data sources, e.g.: for CLABSI, admission data, patient age data (in months, with focus in neonates), laboratory, microbiology, vital signs and imaging reports are necessary to fully comply to the definition. All data sources and their completeness should be checked for deciding the appropriateness of the extracted data for outcome derivation. This is typically done during data preparation. Measurement process - completeness; Representativeness - depth of data - granularity Completeness
Clinical or surveillance definitions are complex and their implementation is prone to variations The definition logic for clinical or surveillance definitions are documented in large documents, typically issued by health authorities like CDC, and the logic is generally complex. As these are replicated in code during the data preparation time using their documentation as the coding requirements, good understanding of the definition is necessary. For example, the VAE definition is time-sensitive and there are multiple corner cases (parts of the definition implementation that require a specific adapted logic). The definitions can be subject of interpretation and there will be variation if different people implement it based on the same documentation. Moreover, coding bugs can occur when implementing complex logic. Unit tests can be foreseen in data preparation for corner cases and time-sensitive interpretation. Good understanding of the definition and its clinical interpretation can be achieved by communication with clinical experts. Whenever other definitions are historically available (e.g.: ICD codes, CLABSI definition manually assessed and documented by infection preventionists), these can be used for detecting deviations. Some deviations might help with correcting programming errors, but not all deviations will necessarily constitute a problem. NA NA
Outcomes derived using clinical or surveillance definition can be affected by missing or erroneous data The retrospective calculation of clinical or surveillance definition relies on the availability of different items in the system, which in turn relies on the hospital process of recording these items. For example, CLABSI or VAE events include fever as a symptom check, which in turn, can be derived by using temperature values. Temperature values can be missing and errors in recording these values will be present. The data recording frequency and its impact on the outcome calculation should be assessed. The necessity of cleaning or imputing the data prior to outcome calculation should be assessed. Alternatively, assumptions can be made (e.g.: microbiology results are always ordered in the presence of fever) to simplify the definition, but such assumptions should be carefully reviewed with clinical experts. Measurement process - completeness; Representativeness - depth of data - granularity Completeness
@@
Specific prediction tasks require the extraction of auxiliary outcomes and/or competing events Deep neural network models may benefit of the use of auxiliary outcomes (multivariate outcomes) as a form of regularization (Tomašev et al., 2021, 2019). Including competing events (like death, discharge, device removal) are sometimes modelled when using regression models (Gao et al., 2024; Thuraisingam et al., 2021). Whenever applicable for the prediction task, the availability of auxiliary outcomes and/or competing event at extraction should be checked. Informativeness - feature importance Completeness
@@
Label leakage due to temporal leakage Whenever the prediction time is too close to the event time, strong features will rather detect than predict the outcome which is already known to the healthcare providers. Such an example is mortality prediction at the disconnection of the mechanical ventilator as a family decision to withdraw care at a terminal stage of the ilness (Ghassemi et al., 2020). Careful definition in the outcome label can alleviate such situations. For example, contact with palliative care, transfer to palliative care or death can be used as outcome definition for mortality prediction. NA NA
{longtabu}
@@
to >p2cm>p5cm>p5cm>p1cm>p1cm FEATURE ENGINEERING Recommendations (not time-sensitive). AVPU = Alert, Voice, Pain, Unresponsive, EHR: Electronic Health Record, ICU: Intensive Care Unit, OMOP CDM: Observational Medical Outcomes Partnership Common Data Mode, ICD: International Classification of Diseases, CCS: Clinical Classifications Software, CCSR: Clinical Classifications Software Refined, ETL: Extract, Transform, Load
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
@@
Inconsistencies due to data extraction of the same clinical items from various databases There will be high variation in data extraction procedure when data sources differ. A typical example are extractions from EHR (hospital-wide) and ICU system of vital signs, devices (e.g.: urinary catheter, dialysis catheter), scales and scores, signs and symptoms or other care observations. One approach is to extract the same items in a harmonized format. Another approach is to extract separate base tables. These often follow different structures (e.g.: units are available or units are implicitly considered the same unit; devices registered with start and end time in ICU, but only through a ’device present’ observation in the EHR), require different mappings, have different significance (e.g.: different pain scales used in different wards or age groups) or can even not be recorded only in one system (e.g.: specific scales are mostly used in ICU, like Glasgow Coma Scale), or recorded in a structured way in one system and in free text in another system (e.g.: ’patient is alert’ in a free text note vs. AVPU scale) The maximum possible information should be extracted from all systems, regardless if data are harmonized during extraction or during data preparation. (e.g.: extract units or start/end date and time if available in one system only). If data sources are harmonized at extraction, include the data source, as cleaning steps might differ for different data sources (e.g.: cleaning temperatures recorded by devices in ICU can be different than cleaning errors in manually entered temperatures in other wards). Standardization at extraction using OMOP CMD helps in harmonizing both the data structure and the clinical concepts terminology across different data sources. Consistency - rule-based consistency - syntactic consistency NA
@@
Same clinical concept available in different tables in the EHR database The registration of specific clinical concepts can deviate from the standard procedure in specialized wards. For example, catheter information is recorded in a standard way for all wards (using nursing notes), but dialysis catheters are recorded in a specialized software module in nephrology department, and implicitly a separate database table. The same can happen with specialized software modules deployed over time, and items being recorded in a database table in older data and in another table in newer data. Additionally, some fields might represent patient reported data while others represent clinically assessed data (Arbet et al., 2021) and depending on the research question, either both or only the clinically assessed data can be relevant. During the clinical concept mapping phase (typically done at data extraction) these recording practices and historical transitions have to be captured accurately from each underlying table. Data exploration during data preparation should reveal only changes in hospital processes if the mapping has been correctly done during extraction. Consistency - rule-based consistency - syntactic consistency NA
@@
Clinical concept mapping differs between wards or over time Even when the same clinical concept is recorded in the same table, configuration catalogues might have changed over time or might differ in different wards. Even small typos (e.g.: veneuze vs. verneuze katheter) can create gaps in data if not mapped correctly. The clinical concept mapping requires knowledge of the historical use of the EHR software in the hospital, including changes in configuration catalogues, and is best achieved through collaboration with IT specialists or clinical experts. Typically this is done at extraction, as a step in the ETL process. OMOP CDM provides standardized mapping, but does not solve the problem of correctly detecting all original items to be mapped. Measurement process - human-introduced error - carelessness; Consistency - rule-based consistency - syntactic consistency Correctness
@@
Partially overlapping data regarding the same clinical concept available in multiple extracted tables Often, multiple extracted tables contain the same clinical information. For example, diagnoses might be coded using ICD codes at the end of the admission, but they might also be available from care pathways recorded during the admission. Alternatively, ICD-9 and ICD-10 coding might coexist (Miao et al., 2023). Medical procedures are coded using CPT codes, typically after discharge, but information on the performed procedures might be available in structured fields in EHR during the patient admission. Medication orders, prescription and administration might represent three different extracted tables. Provided clear extraction documentation, the decision either merging extracted items in one feature or making a choice of a source against the other is generally taken during data preparation. Data source agreement exploration can be done during data preparation. Informativeness - redundancy - conciseness; Consistency - rule-based consistency - compliance; Documentation Concordance
@@
Changes in hospital processes over time result in gaps in the extracted features Gaps in recording features over large periods of time (months or years) can occur. The EHR software or specific software modules are generally deployed ward by ward, and such gaps in data can be seen in the first years of the extracted period, before hospital-wide adoption of the software (e.g.: if emergency department only started using the software since 2014 but data extraction starts in 2012). Gaps can also occur during transitions periods for one coding system to another, e.g.: during the transition period from ICD-9 to ICD-10, the hospital is not legally obliged to code the diagnosis codes for one calendar year. Software bugs occuring for a limited period of time can also create data gaps. Whenever known at extraction, such large impact hospital changes can be documented in the extraction documentation. This warns the data modeller of the usefulness of extracted features for the prediction task. Exploration of trends over time in the recording of each feature (hospital-wide, but also per medical ward) detects these gaps in data. At data preparation time, possible decisions are to not use unreliable features, proceed with missing data imputation or redefine the period of interest or the cohort of interest. Measurement process - completeness; Documentation Completeness
@@
Weekend effects result in missing data Due to staffing constraints, certain clinical items (e.g.: registered by nurses) can be recorded with a lower frequency during weekends or public holidays. Exploring the extracted items or features trends over days of the week can help with detecting such trends. During data preparation, creation of separate features for the days of the week (Monday to Sunday) or weekend flags, eventually together with interactions with the features that display weekend gaps, can help the model distinguish between important patterns during weekdays vs. weekend days. Measurement process - completeness Completeness
@@
Feature set definition is too restrictive Depending on the project approach, data modellers might have to define the set of clinical items to extract at the start of the project. Typically, the important predictors for the outcome of interest are selected, based on literature review and/or collaboration with clinical experts. It can though prove interesting to model the prediction task using a competing events framework (Gao et al., 2024; Albu et al., 2024). Auxiliary outcomes are often used as a form of regularization in deep learning applications (Tomašev et al., 2021). In these situation, extracting items with predictive value for the competing events or auxiliary tasks can prove useful. Additionally, missing data imputation of extracted features can benefit of extraction of more features that have value in imputing the features of interest in the prediction task. Whenever possible, access of the data modeller to the full extraction or extraction of a larger set of clinical items of interest can benefit the data preparation process. Informativeness - feature importance Completeness
@@
Overmapping during extraction can dilute the predictive signal Although terminology mapping during data extraction can reduce significantly the time spent during data preparation, mapping into too broad categories can dilute the predictive signal. For example, ICD to CCS or CCSR mapping might be performed at extraction. Depending on the outcome to be predicted, history of specific ICD codes might prove more predictive than broader categories. Both the more granular extraction as well as the broader mapping could be made available during extraction. Informativeness - feature importance Concordance
@@
Use of hospital aggregate features The prediction task might benefit of the use of aggregate hospital features, such as bed occupancy or nurse/clinician workload at the prediction time, typically aggregated per calendar day. Such features can, in principle, be calculated during data preparation based on the extracted data. At the prediction time in clinical implementation though, the prediction API should receive all data for all patients hospitalized at prediction time for the calculation of patient census or bed occupancy. This constitutes an unnecessarily heavy data transfer for making predictions for one patient at one point in time. Moreover, if the inclusion/exclusion criteria are applied at extraction (e.g.: patients on mechanical ventilation), the researcher does not have access to the precise number of patients that occupy beds in the ward, as not all patients are on mechanical ventilation. It is more convenient to foresee such aggregate features at extraction time (considering that the same logic used at extraction can be reused at clinical implementation prediction time). The same database queries can be easily reused at prediction time, without exposing to the prediction API data for all hospitalised patients. Informativeness - feature importance; Measurement process - completeness Completeness
@@
Use of patient aggregate features Baseline features capturing aggregates of previous patient history, as for example historical comorbidities often prove useful for the prediction task. These can be calculated either at extraction time or at prediction time. Calculating these at data preparation time come though with downsides. First, the extraction might be limited to a specific time range, for example admissions from 2014 to 2020. Patients admitted in 2014 will suffer of left censoring for historical comorbidities, as their admissions prior to 2014 are not included in the extracted data. Second, at clinical implementation time, data for all patient admissions from the past would have to be sent to the prediction API, which might results in unnecessarily large data transfers over the network, especially if predictions are renewed often for a large number of patients. Last, when the exclusion/inclusion criteria are defined at extraction time, not all prior admissions for a patient are available (a patient undergoing surgery during current admission might have had prior admissions without surgery). Although possible (to some extent) to derive such features are data preparation, it might come with limitations. Preferably, these would be calculated at extraction time and identically reproduced in clinical implementation. Within certain limitations, these can also be derived at preparation time. Informativeness - feature importance; Measurement process - completeness Completeness
{longtabu}
@@
to >p2cm>p5cm>p5cm>p1cm>p1cm FEATURE ENGINEERING Recommendations (time-sensitive). EHR: Electronic Health Record, ETL: Extract, Transform, Load, ICD: International Classification of Diseases, CPT: Current Procedural Terminology, AKI: Acute Kidney Injury, CLABSI: Central Line-Associated Bloodstream Infections, ICU: Intensive Care Unit, API: Application Programming Interface
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
Problem Problem description Recommendation METRIC cluster and dimension Weiskopf and Weng dimension
[…]
@@
Timestamps shifting for anonymization As part of the anonymization process applied at data extraction, date shifting can be applied to deidentify the date of patient admissions (Johnson et al., 2016, 2023; Maletzky et al., 2022). This process should preserve the sequence of admissions of a patient, the sequence of events within an admission, as well as seasonal features (e.g.: time of day, day of the week, month of year). However, the process can introduce artifacts and render specific feature extractions at data preparation time impossible (e.g. bed occupancy, nurse workload). The anonymization process applied at extraction should be tested using unit/integration tests in the data extraction framework. Whenever this process renders extraction of certain features unfeasible at data preparation, this should be documented. Timeliness - currency; Documentation Currency

### v2

@@ 4.4 Data cleaning
Outcomes documented in the system are subject to mislabeling and delay in registration DESCRIPTION: Assessments done during the patient’s stay, e.g. delirium assessment by a neurologist using Confusion Assessment Method (CAM) or Delirium Rating Scale (DRS), can be used to define the outcome. These can though be subject of different documentation practices in different wards. For example, delirium might be consistently evaluated on the neurology ward and after surgery, but inconsistently in other wards or in patients that did not undergo surgery. Unless meticulously documented, clinical assessments are subject to inconsistencies in different patient groups or across medical wards. There may also be delays in documentation compared to the clinical assessment time (e.g., documented at the end of the day), which can impact time-sensitive outcomes.
RECOMMENDATION: Understanding the assessment practices of the event(s) identifying the outcome is essential in defining its reliability. Understanding the documentation practices (and the registration delay compared to the actual event time) can also be important, depending on how time-sensitive the predictions are, how often the predictions are renewed and the prediction time horizon. This information is typically retrieved from clinical experts, as documenting every single clinical assessment practice during data extraction would be overwhelming. Manual relabelling by clinical experts post extraction, based on patient assessments, possibly combined with other data such as clinical notes, can prove as a viable solution for some outcomes, although it remains a labour intensive task.
METRIC: Informativeness - understandability; Measurement process - human-introduced error - noisy labels; Measurement process - completeness; Timeliness - currency; WW: Correctness; Currency
@@
Specific prediction tasks require the extraction of auxiliary outcomes and/or competing events DESCRIPTION: Deep neural network models may benefit of the use of auxiliary outcomes (multivariate outcomes, e.g., predicting the levels of creatinine, urea nitrogen, sodium, potassium, chloride, calcium and phosphate, along with AKI) as a form of regularization (Tomašev et al., 2021, 2019). Including competing events (like death, discharge, device removal) are sometimes modelled when using regression models (Gao et al., 2024; Thuraisingam et al., 2021).
RECOMMENDATION: Whenever applicable for the prediction task, the availability of auxiliary outcomes and/or competing event at extraction should be checked.
METRIC: Informativeness - feature importance; WW: Completeness
@@
Label leakage due to temporal leakage DESCRIPTION: Whenever the prediction time is too close to the event time, strong features will rather detect than predict the outcome which is already known to the healthcare providers. Such an example is mortality prediction at the disconnection of the mechanical ventilator as a family decision to withdraw care at a terminal stage of the ilness (Ghassemi et al., 2020). Another example is including timepoints after the occurence of the event of interest, e.g., sepsis, in model training and evaluation. (Wong et al., 2021)
RECOMMENDATION: Careful definition in the outcome label can alleviate such situations. For example, contact with palliative care, transfer to palliative care or death can be used as outcome definition for mortality prediction. The data preparation process should ensure post-event data is excluded when necessary.
METRIC: NA; WW: NA
Table 5: OUTCOME Recommendations for Outcomes Derived Using Clinical or Surveillance definitions. ICD: International Classification of Diseases; AKI: Acute Kidney Injury; CLABSI: Central Line-Associated Bloodstream Infection; VAE: Ventilator-Associated Event; CDC: Centers for Disease Control and Prevention; NA: Not Applicable; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
Outcomes derived using clinical or surveillance definitions often require extraction of multiple data sources DESCRIPTION: Clinical or surveillance definitions provide a more robust alternative compared to ICD codes or documentation of clinical assessments for certain acute events (e.g., AKI, CLABSI, VAE). It comes with advantages: the event occurrence/onset timestamp can be established and the same definition is used over the entire training and test data, but its consistency might still depend on the recording practices of the items used in the definition. Furthermore, these items span across multiple data sources, e.g., for CLABSI, admission data, patient age data (in months, with focus in neonates), laboratory, microbiology, vital signs and imaging reports are necessary to fully comply to the definition.
RECOMMENDATION: All data sources and their completeness should be checked for deciding the appropriateness of the extracted data for outcome derivation. This is typically done during data preparation. Additionally, we recommend extracting both the clinically relevant timestamp as well as the timestamp when data are recorded in the system, whenever these differ (examples in Table 4), and using the clinically relevant timestamp for outcome derivation, e.g. the sample collection time for creatinine when calculating AKI outcome.
METRIC: Measurement process - completeness; Representativeness - depth of data - granularity; WW: Completeness
Clinical or surveillance definitions are complex and their implementation is prone to variations DESCRIPTION: The definition logic for clinical or surveillance definitions are documented in large documents, typically issued by health authorities like CDC, and the logic is generally complex. As these are replicated in code during data preparation using their documentation as the coding requirements, good understanding of the definition is necessary. For example, the VAE definition is time-sensitive and there are multiple corner cases (parts of the definition implementation that require a specific adapted logic). The definitions can be subject of interpretation and there will be variation if different people implement it based on the same documentation. Moreover, coding bugs can occur when implementing complex logic.
RECOMMENDATION: Unit tests can be foreseen in data preparation for corner cases and time-sensitive interpretation. Good understanding of the definition and its clinical interpretation can be achieved by communication with clinical experts. Code reviews can also capture misinterpretations. Whenever other definitions are historically available (e.g., ICD codes, CLABSI definition manually assessed and documented by infection preventionists), these can be used for detecting deviations. Some detected deviations might help with correcting programming errors, but not all deviations will necessarily constitute a problem.
METRIC: NA; WW: NA
Outcomes derived using clinical or surveillance definition can be affected by missing or erroneous data DESCRIPTION: The retrospective calculation of clinical or surveillance definition relies on the availability of different items in the system, which in turn relies on the hospital process of recording these items. For example, CLABSI or VAE events include fever as a symptom check, which in turn, can be derived by using temperature values. Temperature values can be missing and errors in recording these values will be present.
RECOMMENDATION: The data recording frequency and its impact on the outcome calculation should be assessed. The necessity of cleaning or imputing the data prior to outcome calculation should be assessed. Alternatively, assumptions can be made (e.g., microbiology results are always ordered in the presence of fever) to simplify the definition, but such assumptions should be carefully reviewed with clinical experts.
METRIC: Measurement process - completeness; Representativeness - depth of data - granularity; WW: Completeness
@@
Table 6: FEATURE ENGINEERING Recommendations (not time-sensitive) Related to Clinical Concept Mapping. AVPU = Alert, Voice, Pain, Unresponsive; EHR: Electronic Health Record; ICU: Intensive Care Unit; OMOP CDM: Observational Medical Outcomes Partnership Common Data Model; ICD: International Classification of Diseases; ETL: Extract, Transform, Load; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
@@
Inconsistencies due to data extraction of the same clinical items from various databases DESCRIPTION: There will be high variation in data extraction procedure when data sources differ. A typical example are extractions from EHR (hospital-wide) and ICU system of vital signs, devices (e.g., urinary catheter, dialysis catheter), scales and scores, signs and symptoms or other care observations. These often follow different structures (e.g., units, such as cm for neonates length or m for adults height, are available or units are implicitly considered the same unit; devices registered with start and end time in ICU, but only through a ’device present’ observation in the EHR), require different mappings, have different significance (e.g., different pain scales used in different wards or age groups) or can be recorded only in one system (e.g., specific scales are mostly used in ICU, like Glasgow Coma Scale), or recorded in a structured way in one system and in free text in another system (e.g., ’patient is alert’ in a free text note vs. AVPU scale).
RECOMMENDATION: One approach is to extract the same items in a harmonized format. Another approach is to extract separate base tables for each data source. The maximum possible information should be extracted from all systems, regardless if data are harmonized during extraction or during data preparation. (e.g., extract units or start/end date and time if available in one system only). If data sources are harmonized at extraction, include the data source, as cleaning steps might differ for different data sources (e.g., cleaning temperatures recorded by devices in ICU can be different than cleaning errors in manually entered temperatures in other wards). Standardization at extraction using OMOP CMD helps in harmonizing both the data structure and the clinical concepts terminology across different data sources.
Consistency - rule-based consistency - syntactic consistency; WW: NA
@@
Same clinical concept available in different tables in the same database DESCRIPTION: Differences in registration can occur even when data originate in the same database. The registration of specific clinical concepts can deviate from the standard procedure in specialized wards. For example, catheter information may be recorded in a standard structured way for all wards (using nursing observations), but dialysis catheters are recorded in a specialized software module in nephrology department, and implicitly a separate database table. The same can happen with specialized software modules deployed over time, and items being recorded in a database table in older data and in another table in newer data. Additionally, some fields might represent patient reported data while others represent clinically assessed data (Arbet et al., 2021) and depending on the research question, either both or only the clinically assessed data can be relevant.
RECOMMENDATION: During the clinical concept mapping phase (typically done at data extraction) these recording practices and historical transitions have to be captured accurately from each underlying table. Data exploration during data preparation should reveal only changes in hospital processes if the mapping has been correctly done during extraction.
Consistency - rule-based consistency - syntactic consistency; WW: NA
Clinical concept mapping differs between wards or over time DESCRIPTION: Even when the same clinical concept is recorded in the same table, configuration catalogues might have changed over time or might differ in different wards. Configuration changes (e.g., CVC vs. central venous catheter) or even small typos (e.g., centraal vs. central venous catheter) can create gaps in data if not mapped correctly.
@@
RECOMMENDATION: The clinical concept mapping requires knowledge of the historical use of the EHR software in the hospital, including changes in configuration catalogues, and is best achieved through collaboration with IT specialists or clinical experts. Typically, this is done at extraction, as a step in the ETL process. OMOP CDM provides standardized mapping, but does not solve the problem of correctly detecting all original items to be mapped.
Measurement process - human-introduced error - carelessness; Consistency - rule-based consistency - syntactic consistency; WW: Correctness
@@
Partially overlapping data regarding the same clinical concept available in multiple extracted tables DESCRIPTION: Often, multiple extracted tables contain the same clinical information. For example, diagnoses might be coded using ICD codes at the end of the admission, but they might also be available from care pathways recorded during the admission. Alternatively, ICD-9 and ICD-10 coding might coexist (Miao et al., 2023). Medical procedures are coded using CPT codes, typically after discharge, but information on the performed procedures might be available in structured fields in EHR during the patient admission. Medication orders, prescription and administration might represent three different extracted tables.
RECOMMENDATION: Provided clear extraction documentation, the decision either merging extracted items in one feature or making a choice of a source against the other is generally taken during data preparation. Data source agreement exploration can be done during data preparation.
Informativeness - redundancy - conciseness; Consistency - rule-based consistency - compliance; Documentation; WW: Concordance
@@
Changes in hospital processes over time result in gaps in the extracted features DESCRIPTION: Gaps in recording features over large periods of time (months or years) can occur. The EHR software or specific software modules are generally deployed ward by ward, and such gaps in data can be seen in the first years of the extracted period, before hospital-wide adoption of the software (e.g., if emergency department only started using the software since 2014 but data extraction starts in 2012). This, of course, depends on when processes change relative to the extraction period. Gaps can also occur during transitions periods for one coding system to another, e.g., during the transition period from ICD-9 to ICD-10, the hospital is not legally obliged to code the diagnosis codes for one calendar year. Software bugs occurring for a limited period of time can also create data gaps. Recording of some clinical observations may also be discontinued in recent data, for example, to ease the nurses’ workload.
RECOMMENDATION: Whenever known at extraction, such large impact hospital changes can be documented in the extraction documentation. This warns the researcher of the usefulness of extracted features for the prediction task. Exploration of trends over time in the recording of each feature (hospital-wide, but also per medical ward) detects these gaps in data. During data preparation, possible decisions are to not use unreliable features, proceed with missing data imputation or redefine the period of interest or the cohort of interest.
METRIC: Measurement process - completeness; Documentation; WW: Completeness
Table 7: FEATURE ENGINEERING Recommendations (not time-sensitive) Related to Feature Availability and Aggregation. ICD: International Classification of Diseases; CCS: Clinical Classifications Software; CCSR: Clinical Classifications Software Refined; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
Ward-specific data recording patterns DESCRIPTION: Difference between wards will exist in compliance of registration of specific care items and adherence to registration guidelines. These differences may be due to staffing requirements, or may be due to perceived clinical relevance for their specific patient population. Specific scores and scales may be applicable only to specific patient groups. Missingness patterns of extracted items may therefore differ between wards.
RECOMMENDATION: During data preparation, missing data patterns can be explored in each ward and appropriate decisions can be taken depending on context: either impute missing data or engineer features in a contextually relevant manner, e.g. using GCS scores only in appropriate patient groups.
Measurement process - completeness; WW: Completeness
Weekend effects result in missing data DESCRIPTION: Due to staffing constraints, certain clinical items (e.g., registered by nurses) can be recorded with a lower frequency during weekends or public holidays.
@@
RECOMMENDATION: Exploring the extracted items or features trends over days of the week can help with detecting such trends. During data preparation, creation of separate features for the days of the week (Monday to Sunday) or weekend flags, eventually together with interactions with the features that display weekend gaps, can help the model distinguish between important patterns during weekdays vs. weekend days.
Measurement process - completeness; WW: Completeness
@@
Feature set definition is too restrictive DESCRIPTION: Depending on the project approach, researchers might have to define the set of clinical items to extract at the start of the project. Typically, the important predictors for the outcome of interest are selected, based on literature review and/or input from clinical experts. It can though prove interesting to model the prediction task using a competing events framework (Gao et al., 2024; Albu et al., 2024). Auxiliary outcomes are often used as a form of regularization in deep learning applications (Tomašev et al., 2021). In these situations, extracting items with predictive value for the competing events or auxiliary tasks can prove useful. Additionally, missing data imputation of extracted features can benefit of extraction of more features that have value in imputing the features of interest in the prediction task.
RECOMMENDATION: Whenever possible, extraction of a larger set of clinical items of interest can benefit the data preparation process.
METRIC: Informativeness - feature importance; WW: Completeness
@@
Overmapping during extraction can dilute the predictive signal DESCRIPTION: Although terminology mapping during data extraction can reduce significantly the time spent during data preparation, mapping into too broad categories can dilute the predictive signal. For example, ICD to CCS or CCSR mapping might be performed at extraction. Depending on the outcome to be predicted, history of specific ICD codes might prove more predictive than broader categories.
RECOMMENDATION: Both the more granular extraction as well as the broader mapping could be made available during extraction, e.g., both ICD and CCS codes.
Informativeness - feature importance; WW: Concordance
@@
Use of hospital aggregate features DESCRIPTION: The prediction task might benefit of the use of aggregate ward-level features, such as bed occupancy or nurse/clinician workload at the prediction time, typically aggregated per calendar day. Such features can, in principle, be calculated during data preparation based on the extracted data. At the prediction time in clinical implementation though, the prediction API should receive all data for all patients hospitalized at prediction time for the calculation of patient census or bed occupancy. This constitutes an unnecessarily heavy data transfer for making predictions for one patient at one point in time. Moreover, if the inclusion/exclusion criteria are applied at extraction (e.g., patients on mechanical ventilation), the researcher does not have access to the precise number of patients that occupy beds in the ward, as not all patients are on mechanical ventilation.
RECOMMENDATION: It is more convenient to foresee such aggregate features during data extraction, considering that the same logic used at extraction can be reused at clinical implementation prediction time. Reusing the same database queries at prediction time reduces the need for large data transfers and exposure of all hospitalized patient data to the prediction API.
Informativeness - feature importance; Measurement process - completeness; WW: Completeness
@@
Use of patient aggregate features DESCRIPTION: Baseline features capturing aggregates of previous patient history, as for example historical comorbidities often prove useful for the prediction task. These can be calculated either during data extraction or at prediction time. Calculating these during data preparation comes though with downsides. First, the extraction might be limited to a specific time range, for example admissions from 2014 to 2020. Patients admitted in 2014 will suffer of left censoring for historical comorbidities, as their admissions prior to 2014 are not included in the extracted data. Second, at clinical implementation time, data for all patient admissions from the past would have to be sent to the prediction API, which might result in unnecessarily large data transfers over the network, especially if predictions are renewed often for a large number of patients. Last, when the exclusion/inclusion criteria are defined during data extraction, not all prior admissions for a patient are available (a patient undergoing surgery during current admission might have had prior admissions without surgery).
RECOMMENDATION: Preferably, these would be calculated during data extraction and identically reproduced in clinical implementation. Within certain limitations, these can also be derived at preparation time.
Informativeness - feature importance; Measurement process - completeness; WW: Completeness
@@
Table 8: FEATURE ENGINEERING Recommendations (time-sensitive) Related to Timestamp Integrity. EHR: Electronic Health Record; ICD: International Classification of Diseases; AKI: Acute Kidney Injury; CLABSI: Central Line-Associated Bloodstream Infections; ICU: Intensive Care Unit; METRIC = METRIC cluster and dimension mapping; $W$ = Weiskopf and Weng dimension mapping
Problem Description and recommendation
@@
Timestamps shifting for anonymization DESCRIPTION: As part of the anonymization process applied at data extraction, date shifting can be applied to deidentify the date of patient admissions (Johnson et al., 2016, 2023; Maletzky et al., 2022). This process should preserve the sequence of admissions of a patient, the sequence of events within an admission, as well as seasonal features (e.g., time of day, day of the week, month of year). However, the process can introduce artifacts and render specific feature extractions during data preparation impossible (e.g. bed occupancy, nurse workload).
RECOMMENDATION: The anonymization process applied at extraction should be tested using unit/integration tests in the data extraction framework. Whenever this process renders extraction of certain features unfeasible at data preparation, this should be documented.
METRIC: Timeliness - currency; Documentation; WW: Currency